A man spends 70% of his salary on food expenses, 40% of the rest on traveling expenses, 90% of the rest on house rent and was finally left with $3600. How much did he spend on traveling expenses?

Answer by stanbon(75887)

A man spends 70% of his salary on food expenses, 40% of the rest on traveling expenses, 90% of the rest on house rent and was finally left with $3600. How much did he spend on traveling expenses?
--------
Let his salary = "x".
----
Equation:
x - 0.7x - 0.4*0.3x - 0.9[x-0.7x-0.4*3x] = 3600
------
0.6(0.3x) - 0.9[0.6*0.3x)] = 3600
0.1(0.6*0.3x) = 3600
0.018x = 3600
x = 200,000.00 (total salary)
----------------------
Ans: Traveling expenses = 0.4(0.3x) = $24,000.00
---------------------------------------------------
Cheers,
Stan H.
----------------